# Automating Continual Learning

### v1

@@ 3 Method
The ACL objective function above (Eq. 4) is simple but encapsulates desiderata for continual learning (reviewed in Sec. 2.1). The last term of Eq. 4 with $p(y^{\mathcal{A}}_{\text{test}}|{\bm{x}}^{\mathcal{A}}_{\text{test}};{\bm{W}}_{\mathcal{A,B}})$ or schematically ${\bm{p}}(\mathcal{A}^{\prime}|\mathcal{A},\mathcal{B})$, optimizes for backward transfer: (1) remembering the first task ${\mathbf{A}}$ after learning ${\mathbf{B}}$ (combatting catastrophic forgetting), and (2) leveraging learning of ${\mathbf{B}}$ to improve performance on the past task ${\mathbf{A}}$. The second term of Eq. 4, $p(y^{\mathcal{B}}_{\text{test}}|{\bm{x}}^{\mathcal{B}}_{\text{test}};{\bm{W}}_{\mathcal{A,B}})$ or schematically ${\bm{p}}(\mathcal{B}^{\prime}|\mathcal{A},\mathcal{B})$, optimizes forward transfer leveraging the past learning experience of ${\mathbf{A}}$ to improve predictions in the second task ${\mathbf{B}}$, in addition to simply learning to solve Task ${\mathbf{B}}$ from the corresponding training examples. To complete, the first term of Eq. 4 is a simple single-task meta-learning objective for Task ${\mathbf{A}}$.
Overall Model Architecture. As we mention in Sec. 2, in our NN architecture, the core sequential dynamics of continual learning are learned by the self-referential layers. However, as an image-processing NN, our model makes use of a vision backend consisting of multiple convolutional layers. In practice, we use the “Conv-4” architecture of Vinyals et al. (2016) typically used in the context of few-shot learning. Overall, the model takes an image as input, process it through a feedforward convolutional NN, whose output is fed to the SRWM-layer block. Note that this is one of the limitations of this work. While our architecture with fixed vision components still allows us to experimentally demonstrate the principle of ACL, more general ACL should make use of self-modifying NNs that also learn to modify the vision components. One straightforward architecture fitting the bill is an MLP-mixer architecture (Tolstikhin et al. (2021); built of several linear layers), where all linear layers are replaced by the self-referential linear layers of Sec. 2.3. While we implemented such a model, it turned out to be too slow for us to conduct corresponding experiments. Our public code includes a self-referential MLP-mixer implementation, but for further experiments, we leave the future work with such an architecture using more efficient CUDA kernels.
Another crucial architectural choice that is specific to continual/multi-task image processing is normalization layers (see also Bronskill et al. (2020)). Typical NNs used in few-shot learning (e.g., Vinyals et al. (2016)) contain batch normalization (BN; Ioffe & Szegedy (2015)) layers. All our models use instance normalization (IN; Ulyanov et al. (2016)) instead of BN because in our preliminary experiments, as expected, we found IN to generalize much better than BN layers in the CL setting.
4 Experiments
Base Settings. Unless otherwise indicated (e.g, for classic Split-MNIST; Sec. 4.3), all tasks are configured to be a 5-way classification task. This is one of the classic configurations for few-shot learning tasks, and also allows us to evaluate the principle of ACL with reasonable computational costs (like any sequence learning-based meta-learning methods, scaling this to many more classes is challenging; we also discuss this in Sec. 5). For standard datasets such as MNIST, we split the dataset into sub-datasets of disjoint classes (Srivastava et al., 2013): for example for MNIST which is originally a 10-way classification task, we split it into two 5-way tasks, one consisting of images of class ‘0’ to ‘4’ (‘MNIST-04’), and another one made of class ‘5’ to ‘9’ images (‘MNIST-59’). When we refer to a dataset without specifying the class range, we refer to the first sub-set. Unless stated otherwise, we concatenate 15 examples for each class for each task in the context for both meta-training and meta-testing (resulting in sequences of length 75 for each task). All images are resized to $32\times 32$-size $3$-channel images, and normalized according to the original dataset statistics. Appendix A provides further details.
@@ 4 Experiments
Table 1: 5-way classification accuracies using 15 (meta-test training) examples for each class in the context. Each row corresponds to a single model. Bold numbers highlight cases where in-context catastrophic forgetting is avoided through ACL.
@@
4.1 Two-Task Evaluation
We first demonstrate the essence of our method (Sec. 3) in the two-task “domain-incremental” CL setting (see Appendix A.1). We consider two meta-training task combinations: Omniglot (Lake et al., 2015) and Mini-ImageNet (Vinyals et al., 2016; Ravi & Larochelle, 2017) or FC100 (Oreshkin et al., 2018) (which is based on CIFAR100 (Krizhevsky, 2009)) and Mini-ImageNet. The order of appearance of two tasks within meta-training sequences is alternated for every batch. We compare systems trained with or without the backward transfer term in the ACL loss (the last term in Eq. 4).
Table 2 shows the results when the models are meta-tested on the test sets of the corresponding few-shot learning datasets used for meta-training. We observe that in both pairs of training tasks, the models without the ACL loss catastrophically forget the first task after learning the second one: the accuracy on the first task is at the chance level of about 20% for 5-way classification after learning the second task in-context. The ACL loss clearly addresses this problem: the ACL-learned CL algorithms preserve the performance of the first task. This effect is particularly pronounced in the Omniglot/Mini-ImageNet case (involving two very different domains).
Table 2 shows a similar evaluation but on two standard datasets, 5-way MNIST and CIFAR-10. Again, ACL-trained models better preserve the memory of the first task after learning the second one. In the Omniglot/Mini-ImageNet case, we even observe certain positive backward tranfer effects: in particular, in the “MNIST-then-CIFAR10” continual learning case, the performance on MNIST noticeably improves after learning CIFAR10 (possibly leveraging ‘more data’ provided in-context).
4.2 In-Context Catastrophic Forgetting in the Baselines
Here we report some empirical observations on the baseline models trained without the backward transfer term (the last/third term in Eq. 4) in the ACL objective loss (corresponding to the ACL/No cases in Tables 2 and 2), namely emergence of “in-context catastrophic forgetting” during training. We focus on the Omniglot/Mini-ImageNet case, but similar trends can also be observed in the FC100/Mini-ImageNet case. Figures 2a and 2b show two cases we typically observe. These figures show an evolution of six individual meta-training loss terms (the lower the better), reported separately for the cases where Task A (here Omniglot) or Task B (here Mini-ImageNet) appears at the first (1) or second (2) position in the 2-task continual learning training sequences. 4 out of 6 curves correspond to the learning progress, showing whether the model becomes capable of in-context learning the given task (A or B) at the given position (1 or 2). The 2 remaining curves are the ACL backward tranfer loss, also measured for Task A and B separately here.
Figure 2a shows the case where two tasks are learned about at the same time. We observe that when the learning curves go down, the ACL losses go up, indicating that more the model learns, more it tends to forget the task in-context learned previously. This trend is similar when one task is learned before the other one as is the case in Figure 2b. Here Task A alone is learned first; while Task B is not learned, both learning and ACL curves go down for Task A (essentially, as the model does not learn the second task, there is no force that encourages forgetting). At around 3000 steps, the model also starts learning Task B. From this point, the ACL loss for Task A also starts to go up, indicating again an opposing force effect between learning a new task and remembering a past task. These observations clearly indicate that, without explicitly taking into account the backward transfer loss as part of learning objectives, the gradient descent search tends to find solutions/CL algorithms that prefer to erase previously learned knowledge (this is rather intuitive; it seems easier to find such algorithms that ignore any influence of the current learning to past learning than those that also preserve prior knowledge). In all cases, we find our ACL objective to be crucial for the learned CL algorithms to be capable of remembering the old task while also learning the new task at the same.
@@ 4.2 In-Context Catastrophic Forgetting in the Baselines
4.3 General Evaluation
Table 3: Classification accuracies (%) on the Split-MNIST domain-incremental (DIL) and class-incremental learning (CIL) settings (Hsu et al., 2018). Both tasks are 5-task CL problems. For the CIL case, we also report the 2-task case for which we can directly evaluate our out-of-the-box ACL meta-learner of Sec. 4.1 (trained with a 5-way output and the 2-task ACL loss) which, however, is not applicable (N.A.) to the 5-task CIL requiring a 10-way output. Mean/std over 10 training/meta-testing runs. None of the methods here requires replay memory. See Appendix A.7/B.1 for further details.
@@ 4.3 General Evaluation
Online-aware Meta Learning (OML; Out-of-the-box) 69.9 $\pm$ 2.8 46.6 $\pm$ 7.2 24.9 $\pm$ 4.1
@@
 + meta-finetuned with 5-task ACL loss, Omniglot 84.3 $\pm$ 1.2 93.4 $\pm$ 1.2 74.6 $\pm$ 2.3
@@
Evaluation on standard Split-MNIST. Here we evaluate ACL on the standard Split-MNIST task in domain-incremental and class-incremental settings (Hsu et al., 2018; Van de Ven & Tolias, 2018b), and compare its performance to existing CL and meta-CL algorithms (see Appendix A.7 for full references of these methods). Our comparison focuses on methods that do not require replay memory. Table 3 shows the results. Since our ACL-trained models are general-purpose learners, they can be directly evaluated (meta-tested) on a new task, here Split-MNIST. “ACL (Out-of-the-box model)” row of Table 3 corresponds to our model from Sec. 4.1 meta-trained on Omniglot and Mini-ImageNet using the 2-task ACL objective. It performs very competitively with the best existing methods in the domain-incremental setting, while it largely outperforms them in the 2-task class-incremental setting. The same model can be further meta-finetuned using the 5-task version of the ACL loss (here we only used Omniglot as the meta-training data). The resulting model (the last row of Table 3) outperforms all other methods in all settings studied here. We refer to Appendix A.7/B.1 for further discussions.
@@
Evaluation on diverse task domains. Here we evaluate our ACL-trained models for CL involving more tasks/domains; using meta-test sequences made of MNIST, CIFAR-10, and Fashion MNIST (FMNIST). We also evaluate the impact of the number of tasks in the ACL objective: in addition to the model meta-trained on Omniglot/Mini-ImageNet (Sec. 4.1), we also meta-train a model (with the same architecture and hyper-parameters) using 3 tasks, Omniglot, Mini-ImageNet, and FC100, using the 3-task ACL objective (see Appendix A.5); which is not only meta-trained for longer CL, but also uses more data. Table 4 shows the results. The two ACL-trained models are indeed capable of retaining the knowledge without catastrophic forgetting for multiple tasks during meta-testing, while the performance on prior tasks gradually degrades as the model learns new tasks, and performance on new tasks becomes moderate (see Sec. 5 on limitations). The 3-task version outperforms the 2-task one overall, encouragingly indicating a potential for further improvements even with a fixed parameter count.
@@
Table 4: 5-way classification accuracies using 15 examples for each class for each task in the context. 2-task models are meta-trained on Omniglot and Mini-ImageNet, while 3-task models are in addition meta-trained on FC100. ‘A, B’ in ‘Context/Train’ column indicates that models sequentially observe meta-test training examples of Task A then B; evaluation is only done at the end of the sequence.
@@ 5 Discussion
Prior work. While there are several recent works that are catagorized as ‘meta-continual learning’ or ‘continual meta-learning’ (see, e.g., Javed & White (2019); Beaulieu et al. (2020); Caccia et al. (2020); He et al. (2019); Yap et al. (2021); Munkhdalai & Yu (2017), most of them are based on “model-agnostic meta-learning” (MAML; Finn et al. (2017); Finn & Levine (2018)) and learn representations for CL but still make use of classic CL algorithms. In particular, tuning of the learning rate and number of iterations is still required for optimal performance (see, e.g., Appendix A.7). In contrast, our approach learn learning algorithms in the spirit of Hochreiter et al. (2001); Younger et al. (1999); this may be categorized as ‘in-context continual learning.’ Several recent works (see, e.g., Irie & Schmidhuber (2023a); Coda-Forno et al. (2023); von Oswald et al. (2023b)) mention the possibility of such in-context CL without any concrete study. We show that in-context learning also suffers from catastrophic forgetting (Sec. 4.1-4.2) and propose ACL to address this problem. We also note that the use of SRWM is relevant to ‘continual meta-learning’ since with a regular sequence processor with slow weights, there remains the question of how to continually learn the slow weights (meta-parameters). In principle, recursive self-modification as in SRWM is an answer to this question as it collapses such meta-levels into single self-reference (Schmidhuber, 1992a). We also refer to Schmidhuber (1994; 1995); Schmidhuber et al. (1997) for other prior work on meta-continual learning.
@@
Artificial v. Natural ACL in Large Language Models? Recently, “on-the-fly” few-shot/meta learning capability of sequence processing NNs has attracted broader interests in the context of large language models (LLMs; Radford et al. (2019)). In fact, the task of language modeling itself has a form of sequence processing with error feedback (essential for meta-learning (Schmidhuber, 1990)): the correct label to be predicted is fed to the model with a delay of one time step in an auto-regressive manner. Trained on a large amount of text covering a wide variety of credit assignment paths, LLMs exhibit certain sequential few-shot learning capabilities in practice (Brown et al., 2020). This was rebranded as in-context learning, and has been the subject of numerous recent studies (e.g., Xie et al. (2022); Min et al. (2022); Yoo et al. (2022); Chan et al. (2022b; a); Kirsch et al. (2022); Akyürek et al. (2023); von Oswald et al. (2023a); Dai et al. (2023)). Here we explicitly/artificially construct ACL meta-training sequences and objectives, but in modern LLMs trained on a large amount of data mixing a large diversity of dependencies using a large backpropagation span, it is conceivable that some ACL-like objectives may naturally appear in the data.
@@
Limitations. While we argue that our ACL is a promising approach for automating development of CL algorithms, there are also several limitations. First of all, directly scaling ACL for real-world tasks requiring many more classes does not seem straightforward: it would require very long training sequences. That said, it is also possible that ACL could be achieved without exactly following the process we propose; as we mention above for the case of LLMs, certain real-world data may naturally give rise to an ACL-like objective. Another limitation of this work is training within the limited span. We noted that unlike the standard “quadratic” Transformers, linear Transformers/FWPs-like SRWMs can be trained by carrying over states across two consecutive batches for arbitrarily long sequences. Such an approach has been successfully applied to language modeling with FWPs (Schlag et al., 2021). This possibility, however, has not been investigated here, and is left for future work. This work is also limited to the task of image classification, which can be solved by feedforward NNs. Future work may investigate the possibility to extend ACL to continual learning of sequence learning tasks, such as continually learning new languages. Finally, ACL learns CL algorithms that are specific to the pre-specified model architecture; more general meta-learning algorithms may aim at achieving learning algorithms that are applicable to any model, as is the case for many classic learning algorithms.

### v2

@@ 3 Method
The ACL objective function above (Eq. 4) is simple but encapsulates desiderata for continual learning (Sec. 2.1). The last term of Eq. 4 with $p(y^{\mathcal{A}}_{\text{test}}|{\bm{x}}^{\mathcal{A}}_{\text{test}};{\bm{W}}_{\mathcal{A,B}})$ or schematically ${\bm{p}}(\mathcal{A}^{\prime}|\mathcal{A},\mathcal{B})$, optimizes for backward transfer: (1) remembering the first task ${\mathbf{A}}$ after learning ${\mathbf{B}}$ (combatting catastrophic forgetting), and (2) leveraging learning of ${\mathbf{B}}$ to improve performance on the past task ${\mathbf{A}}$. The second term of Eq. 4, $p(y^{\mathcal{B}}_{\text{test}}|{\bm{x}}^{\mathcal{B}}_{\text{test}};{\bm{W}}_{\mathcal{A,B}})$ or schematically ${\bm{p}}(\mathcal{B}^{\prime}|\mathcal{A},\mathcal{B})$, optimizes forward transfer leveraging the past learning experience of ${\mathbf{A}}$ to improve predictions in the second task ${\mathbf{B}}$, in addition to simply learning to solve Task ${\mathbf{B}}$ from the corresponding training examples. To complete, the first term of Eq. 4 is the single-task meta-learning objective for Task ${\mathbf{A}}$.
@@
Table 1: 5-way classification accuracies using 15 (meta-test training) examples for each class in the context. Each row is a single model. Bold numbers highlight cases where in-context catastrophic forgetting is avoided through ACL.
@@
Overall Model Architecture. As we mention in Sec. 2, in our NN architecture, the core sequential dynamics of CL are learned by the self-referential layers. However, as an image-processing NN, our model makes use of a vision backend consisting of multiple convolutional layers. In practice, we use the “Conv-4” architecture of Vinyals et al. (2016) typically used in the context of few-shot learning. Overall, the model takes an image as input, process it through a feedforward convolutional NN, whose output is fed to the SRWM-layer block. Note that this is one of the limitations of this work. While our architecture with fixed vision components still allows us to demonstrate the principle of ACL, more general ACL should make use of self-modifying NNs that also learn to modify the vision components. [^1] One straightforward architecture fitting the bill is an MLP-mixer architecture (Tolstikhin et al. (2021); built of several linear layers), where all linear layers are replaced by the self-referential linear layers of Sec. 2.3. While we implemented such a model, it turned out to be too slow for us to conduct corresponding experiments. Our public code will include a “self-referential MLP-mixer” implementation, but for further experiments, we leave the future work on such an architecture using more efficient CUDA kernels.
Another crucial architectural choice that is specific to continual/multi-task image processing is normalization layers (see also Bronskill et al. (2020)). Typical NNs used in few-shot learning (e.g., Vinyals et al. (2016)) contain batch normalization (BN; Ioffe & Szegedy (2015)) layers. All our models use instance normalization (IN; Ulyanov et al. (2016)) instead of BN because in our preliminary experiments, we expectably found IN to generalize much better than BN layers in the CL setting.
@@
Table 3: Classification accuracies (%) on the Split-MNIST domain-incremental (DIL) and class-incremental learning (CIL) settings (Hsu et al., 2018). Both tasks are 5-task CL problems. For the CIL case, we also report the 2-task case for which we can directly evaluate our out-of-the-box ACL meta-learner of Sec. 4.1 (trained with a 5-way output and the 2-task ACL loss) which, however, is not applicable (N.A.) to the 5-task CIL requiring a 10-way output. Mean/std over 10 training/meta-testing runs. None of the methods here requires replay memory. See Appendix A.7 & B for further details and discussions.
@@
Online-aware Meta Learning (OML) 69.9 $\pm$ 2.8 46.6 $\pm$ 7.2 24.9 $\pm$ 4.1
@@
+ meta-finetuned with 5-task ACL loss, Omniglot 84.5 $\pm$ 1.6 96.0 $\pm$ 1.0 84.3 $\pm$ 1.2
4 Experiments
Common Settings. Unless otherwise indicated (e.g, for classic Split-MNIST; Sec. 4.3), all tasks are configured to be a 5-way classification task. This is one of the classic configurations for few-shot learning tasks, and also allows us to evaluate the principle of ACL with reasonable computational costs (like any sequence learning-based meta-learning methods, scaling this to many more classes is challenging; we also discuss this in Sec. 5). For standard datasets such as MNIST, we split the dataset into sub-datasets of disjoint classes (Srivastava et al., 2013): for example for MNIST which is originally a 10-way classification task, we split it into two 5-way tasks, one consisting of images of class ‘0’ to ‘4’ (‘MNIST-04’), and another one made of class ‘5’ to ‘9’ images (‘MNIST-59’). When we refer to a dataset without specifying the class range, we refer to the first sub-set. Unless stated otherwise, we concatenate 15 examples from each class for each task in the context for both meta-training and meta-testing (resulting in sequences of length 75 for each task). All images are resized to $32\times 32$-size $3$-channel images, and normalized according to the original dataset statistics. We refer to Appendix A for further details.
4.1 Two-Task Evaluation
We first demonstrate the essence of our method (Sec. 3) in the two-task “domain-incremental” CL setting (see Appendix A.1). We consider two meta-training task combinations: Omniglot (Lake et al., 2015) and Mini-ImageNet (Vinyals et al., 2016; Ravi & Larochelle, 2017) or FC100 (Oreshkin et al., 2018) (which is based on CIFAR100 (Krizhevsky, 2009)) and Mini-ImageNet. The order of appearance of two tasks within meta-training sequences is alternated for every batch. Appendix A.2 provides further details. We compare systems trained with or without the backward transfer term in the ACL loss (the last term in Eq. 4).
Table 2 shows the results when the models are meta-tested on the test sets of the corresponding few-shot learning datasets used for meta-training. We observe that for both pairs of meta-training tasks, the models without the ACL loss catastrophically forget the first task after learning the second one: the accuracy on the first task is at the chance level of about 20% for 5-way classification after learning the second task in-context (see rows with “ACL No”). The ACL loss clearly addresses this problem: the ACL-learned CL algorithms preserve the performance of the first task. This effect is particularly pronounced in the Omniglot/Mini-ImageNet case (involving two very different domains).
Table 2 shows evaluations of the same models but using two standard datasets, 5-way MNIST and CIFAR-10, for meta-testing. Again, ACL-trained models better preserve the memory of the first task after learning the second one. In the Omniglot/Mini-ImageNet case, we even observe certain positive backward tranfer effects: in particular, in the “MNIST-then-CIFAR10” continual learning case, the performance on MNIST noticeably improves after learning CIFAR10 (possibly leveraging ‘more data’ provided in-context).
4.2 In-Context Catastrophic Forgetting in the Baselines
Now we closely look at the emergence of “in-context catastrophic forgetting” during meta-training for the baseline models trained without the backward transfer term (the last/third term in Eq. 4) in the ACL objective loss (corresponding to the ACL/No cases in Tables 2 and 2). We focus on the Omniglot/Mini-ImageNet case, but similar trends can also be observed in the FC100/Mini-ImageNet case. Figures 2a and 2b show two representative cases we typically observe. These figures show an evolution of six individual meta-training loss terms (the lower the better), reported separately for the cases where Task A (here Omniglot) or Task B (here Mini-ImageNet) appears at the first (1) or second (2) position in the 2-task CL meta-training training sequences. 4 out of 6 curves correspond to the learning progress, showing whether the model becomes capable of in-context learning the given task (A or B) at the given position (1 or 2). The 2 remaining curves are the ACL backward tranfer losses, also measured for Task A and B separately here.
Figure 2a shows the case where two tasks are learned about at the same time. We observe that when the learning curves go down, the ACL losses go up, indicating that more the model learns, more it tends to forget the task in-context learned previously. We also find this same trend when one task is learned before the other one as is the case in Figure 2b. Here Task A alone is learned first; while Task B is not learned, both learning and ACL curves go down for Task A (essentially, as the model does not learn the second task, there is no force that encourages forgetting). After around 3000 steps, the model also starts learning Task B. From this point, the ACL loss for Task A also starts to go up, indicating again an opposing force effect between learning a new task and remembering a past task. These observations clearly indicate that, without explicitly taking into account the backward transfer loss as part of learning objectives, our gradient descent search tends to find solutions/CL algorithms that prefer to erase previously learned knowledge (this is rather intuitive; it seems easier to find such algorithms that ignore any influence of the current learning to past learning than those that also preserve prior knowledge). In all cases, we find our ACL objective to be crucial for the learned CL algorithms to be capable of remembering the old task while also learning the new task at the same.
4.3 General Evaluation
@@
Evaluation on standard Split-MNIST. Here we evaluate ACL on the standard Split-MNIST task in domain-incremental and class-incremental settings (Hsu et al., 2018; Van de Ven & Tolias, 2018b), and compare its performance to existing CL and meta-CL algorithms (see Appendix A.7 for full references of these methods). Our comparison focuses on methods that do not require replay memory. Table 3 shows the results. Since our ACL-trained models are general-purpose learners, they can be directly evaluated (meta-tested) on a new task, here Split-MNIST. The second-to-last row of Table 3, “ACL (Out-of-the-box model)”, corresponds to our model from Sec. 4.1 meta-trained on Omniglot and Mini-ImageNet using the 2-task ACL objective. It performs very competitively with the best existing methods in the domain-incremental setting, while it largely outperforms them (all but another meta-CL method, GeMCL) in the 2-task class-incremental setting. The same model can be further meta-finetuned using the 5-task version of the ACL loss (here we only used Omniglot as the meta-training data). The resulting model (the last row of Table 3) outperforms all other methods in all settings studied here. We are not aware of any existing hand-crafted CL algorithms that can achieve ACL’s performance without any replay memory. We refer to Appendix A.7/B for further discussions and ablation studies.
@@
Evaluation on diverse task domains. Here we evaluate our ACL-trained models for CL involving more tasks/domains; using meta-test sequences made of MNIST, CIFAR-10, and Fashion MNIST (FMNIST). We also evaluate the impact of the number of tasks in the ACL objective: in addition to the model meta-trained on Omniglot/Mini-ImageNet (Sec. 4.1), we also meta-train a model (with the same architecture and hyper-parameters) using 3 tasks, Omniglot, Mini-ImageNet, and FC100, using the 3-task ACL objective (see Appendix A.5); which is meta-trained not only on longer CL sequences but also on more data. Table 4 shows the results. The two ACL-trained models are indeed capable of retaining the knowledge without catastrophic forgetting for multiple tasks during meta-testing, while the performance on prior tasks gradually degrades as the model learns new tasks, and performance on new tasks becomes moderate (see also Sec. 5 on limitations). The 3-task version outperforms the 2-task one overall, encouragingly indicating a potential for further improvements even with a fixed parameter count.
@@
Table 4: 5-way classification accuracies using 15 examples for each class for each task in the context. 2-task models are meta-trained on Omniglot and Mini-ImageNet, while 3-task models are in addition meta-trained on FC100. ‘A, B’ in ‘Context/Train’ column indicates that models sequentially observe meta-test training examples of Task A then B; evaluation is only done at the end of the sequence. “no ACL” is the baseline 2-task models trained without the ACL loss.
@@ 5 Discussion
Related work. There are several recent works that are catagorized as ‘meta-continual learning’ or ‘continual meta-learning’ (see, e.g., Javed & White (2019); Beaulieu et al. (2020); Caccia et al. (2020); He et al. (2019); Yap et al. (2021); Munkhdalai & Yu (2017)). For example, Javed & White (2019); Beaulieu et al. (2020) use “model-agnostic meta-learning” (MAML; Finn et al. (2017); Finn & Levine (2018)) to meta-learn representations for CL while still making use of classic learning algorithms for CL; this requires tuning of the learning rate and number of iterations for optimal performance during CL at meta-test time (see, e.g., Appendix A.7). In contrast, our approach learn learning algorithms in the spirit of Hochreiter et al. (2001); Younger et al. (1999); this may be categorized as ‘in-context continual learning.’ Several recent works (see, e.g., Irie & Schmidhuber (2023a); von Oswald et al. (2023b)) mention the possibility of such in-context CL but existing works (Irie et al., 2022c; Coda-Forno et al., 2023; Lee et al., 2023) that learn multiple tasks sequentially in-context do not focus on catastrophic forgetting which is one of the central challenges of CL. Here we show that in-context learning also suffers from catastrophic forgetting in general (Sec. 4.1-4.2) and propose ACL to address this problem. We also note that the use of SRWM is relevant to ‘continual meta-learning’ since with a regular sequence processor with slow weights, there remains the question of how to continually learn the slow weights (meta-parameters). In principle, recursive self-modification as in SRWM is an answer to this question as it collapses such meta-levels into single self-reference (Schmidhuber, 1992a). We also refer to Schmidhuber (1994, 1995); Schmidhuber et al. (1997) for other prior work on meta-continual learning.
@@
Artificial v. Natural ACL in Large Language Models? Recently, “on-the-fly” few-shot/meta learning capability of sequence processing NNs has attracted broader interests in the context of large language models (LLMs; Radford et al. (2019)). In fact, the task of language modeling itself has a form of sequence processing with error feedback (essential for meta-learning (Schmidhuber, 1990)): the correct label to be predicted is fed to the model with a delay of one time step in an auto-regressive manner. Trained on a large amount of text covering a wide variety of credit assignment paths, LLMs exhibit certain sequential few-shot learning capabilities in practice (Brown et al., 2020). This was rebranded as in-context learning, and has been the subject of numerous recent studies (e.g., Xie et al. (2022); Min et al. (2022); Yoo et al. (2022); Chan et al. (2022b, a); Kirsch et al. (2022); Akyürek et al. (2023); von Oswald et al. (2023a); Dai et al. (2023)). Here we explicitly/artificially construct ACL meta-training sequences and objectives, but in modern LLMs trained on a large amount of data mixing a large diversity of dependencies using a large backpropagation span, it is conceivable that some ACL-like objectives may naturally appear in the data.
@@
Limitations. While we present our ACL as a promising approach for automating development of CL algorithms, there are also several limitations. First of all, as an in-context/learned learning algorithm, there are challenges in terms of both domain and length generalization (we qualitatively observe these to some extent in Sec. 4; further discussion and experimental results are presented in Appendix B.3 & B.4). Regarding the length generalization, we note that unlike the standard “quadratic” Transformers, linear Transformers/FWPs-like SRWMs can be trained by carrying over states across two consecutive batches for arbitrarily long sequences. Such an approach has been successfully applied to language modeling with FWPs (Schlag et al., 2021). This possibility, however, has not been investigated here, and is left for future work. Also, directly scaling ACL for real-world tasks requiring many more classes does not seem straightforward: it would require very long training sequences. That said, it may be possible that ACL could be achieved without exactly following the process we propose; as we mention above for the case of LLMs, certain real-world data may naturally give rise to an ACL-like objective. This work is also limited to the task of image classification, which can be solved by feedforward NNs. Future work may investigate the possibility to extend ACL to continual learning of sequence learning tasks, such as continually learning new languages. Finally, ACL learns CL algorithms that are specific to the pre-specified model architecture; more general meta-learning algorithms may aim at achieving learning algorithms that are applicable to any model, as is the case for many classic learning algorithms.